a dog weighing 20 kg is in a wagon moving at 3 m/sec on a frictionless surface. a 10 kg pet monkey jumps straight down from a ledge onto the. dogs back for a free ride. what is the velocity of the duo after the monkey lands on the dogs back
Answer #1

initial momentum = m*v = 20*3 = 60 kgm/sec

after monkey jumps total mass = 20 + 10 = 30

Initial momentum = final momentum

60 = 30*v

v = 2 m/s